OpenStudy (anonymous):

Find the volume of a cone with a diameter of 7 inches and a height that is 4 times the radius. Use 3.14 for pi, and round your answer to the nearest hundredth. 29.31 in3 87.92 in3 117.23 in3 179.50 in3

OpenStudy (igreen):

\(V = \dfrac{1}{3} \pi r^2 h\)

OpenStudy (igreen):

Diameter = 7 Radius = 3.5 Height = 14

OpenStudy (igreen):

\(V = \dfrac{1}{3}(3.14)(3.5^2)(14)\) Simplify

OpenStudy (anonymous):

179

OpenStudy (igreen):

Yep, that's our answer..

OpenStudy (anonymous):

its d